Shift Summary for owl shift, 07/01/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Thu, 1 Jul 1999 07:01:23 -0500

Shift Summary for owl shift, 07/01/99

On Shift: Sandrine, Leo

Hours of beam delivered at all intensities: 6
Hours of beam delivered at nominal intensity: 6
Hours of E832 data (special runs not included): 6
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

- around 30 min. lost because the beam stop was disabled
- Tevatron seemed to have some problems (magnet f17lam tripps off, rad.
trip. booster...), beam went off and back many many times after 3am.

========================================================================
Runs ended this shift:
Run 13777 - 12.8 - BEAM_832 still continuing.

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

good e832 data taken for almost the entire shift.
No major problems.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13777 0:00 Inherit shift. Muon run just ended; 4 channels
at the low end of warning range. 832 run starts.

13777 ? 0:05 Sandrine has dinner. Leo's stomach starts to
rumble.
Pasta. His favorite foodstuff.

13777 34 1:02 KTEVANA errors concerning top of stream 5 ->
kfrend.

13777 73 1:47 Channel 689 has a laser stability RMS=3.5%
1231 " 4.6%
but this is still below the panic level.

13777 102 2:26 Little out of limit of intrinsic DC efficiency,
but only at 1 sigma level.

13777 3:00 Second little out-of-limit (when a shift is so
quiet,
you desesperate to find such a "little out of
limit"..)
in the ADC of a channel of V1 plane.

13777 127 3:10 Scaler for CA is out range ...It turns out
that the beam stop is disabled for a short time.
MCR notified.

13777 151 3:40 Beam stop enabled, everything back to normal.
Spills 127 to 151 are bad spills.

3:50 Leo has dinner. It smells good, has bones, even
probably had wings some days ago...

13777 3:57 Explanation of the previous trip: magnet f17lam of
the Tevatron had tripped off. They seem to have
some recurrent problems up there. Beam goes off
and back every 5 mins.

4:15 Gas check list completed.

13777 4:30 MU2 ADC mean profile: channel 28 is upper the
limit.
The HV should be lowered in the next access.

13777 202 4:44 KTEVANA error concerning stream 5. KFREND fixed it.

13777 5:40 Same problem, same solution.

13777 295 6:45 Saw red light on SDAQ pannel for failed +6V on
Station 5, rack 1, lower CAMAC crate. This rack
has DC3, SCIA; both histos look OK and the error
soon disappears.

========================================================================
Follow up Notes on Previous Shift

LASER SCAN not done, since we've never down.

========================================================================
Items to Follow up on Next Shift

LASER SCAN still needs to be done.